OpenStudy (magnetcross):

Can someone help me please? Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y − 4 = 3(x − 2) y − 2 = 3(x − 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

OpenStudy (magnetcross):

will give metal

OpenStudy (welshfella):

PLease don't give direct answers Read the Code of Conduct.

OpenStudy (welshfella):

these equations are given in point slope form which is y - y1 = m(x - x1) where m = slope and the point on the line is (x1,y1)

OpenStudy (welshfella):

You are given the slope which is 3 and the point (x1,y1) is (2,4) so all you need do in replace x1, y1 and m in the above equation with these values